# Weight Space Representation Learning with Neural Fields

### v1

@@ 3.5 Diffusion Model on Weight Representations
We parameterize the reverse process using a diffusion transformer (DiT) peebles2022scalable that predicts the noise added to the weights. For standalone MLP weights, we adopt the architecture from erkocc2023hyperdiffusion; peebles2023learning. For LoRA weights, we design a hierarchical LoRA layer encoder module that respects the structural properties of low-rank weight matrices, shown in Figure 2. Each layer $l$ is processed as follows. First, we treat each vector pair $(\mathbf{a}_{l}^{(i)},\mathbf{b}_{l}^{(i)})$ as a token, where $\mathbf{a}_{l}^{(i)}$ and $\mathbf{b}_{l}^{(i)}$ are the $i$-th columns and rows of matrices $\mathbf{A}^{(l)}$ and $\mathbf{B}^{(l)}$, respectively. Vector-level positional encodings are then applied to capture the rank dimension index. A multi-head attention module with $r$ attention heads encodes the interactions among the $r$ vector pairs within the layer, allowing the model to learn dependencies between different rank components. Finally, layer-level positional encodings are applied to the aggregated layer representation before feeding into the main transformer.
@@ 4 Experiments
Datasets. We evaluate on two datasets. For 2D data, we use FFHQ karras2019style, which contains high quality face images. We evaluate at the resolution of $128\times 128$, although modest compared to current state-of-the-art in image generation, this resolution is significantly higher than what previous weight-space methods zhou2023nft; zhou2023permutation have evaluated on and therefore more challenging. For 3D data, we use ShapeNet chang2015shapenet, focusing on two settings: A single-category model trained on airplanes, and a multi-category model trained on ten object categories including airplanes, chairs, tables, and other common objects.
@@ 4.1 Reconstruction
Results. As shown in Table 1, LoRA, mLoRA and mLoRA-Asym achieve better reconstruction quality while maintaining a compact parameter count, compared to the MLP-based weight representations. We attribute this to the inductive bias provided by the base network, which captures transferable features across instances, and fine tuning effectively leverages these shared representations with minimal adaptation parameters. Multiplicative LoRA perform better than its additive counterpart, echoing the wide application of multiplicative modulation in generative neural fields anokhin2021cips; chan2020pigan; karras2021alias. Surprisingly, mLoRA-Asym consistently outperforms mLoRA despite having certain parameters frozen by the asymmetric mask. Our hypothesis is that the masking further reduces parameter entanglement and thereby improves the reconstruction accuracy. Also notably, LoRA-Asym performs poorly on FFHQ, likely due to increased parameter entanglement caused by large variance $\kappa=6$ (empirically determined) of the frozen weights, as discussed in Section  3.4.
@@ 4.2 Weight Space Structure Analysis
Second, to examine linear mode connectivity frankle2020linear, we measure the barrier height by evaluating reconstruction quality at the midpoint of the linear interpolation path. Specifically, we compute the Chamfer Distance between the ground truth mesh vertices $\mathbf{v}_{\text{gt}}$ and the vertices $\mathbf{v}_{\text{avg}}$ extracted via marching cubes from the averaged weights, i.e.,
@@
Although not always improving reconstruction quality, the asymmetric mask improves both weight similarity and linear mode connectivity across all parameterizations. Notably, mLoRA-Asym exhibits exceptional behavior: weight similarity remains very high and the barrier remains very low even with very different initializations. This suggests that mLoRA-Asym weights converge to a linear mode. The reason is likely that multiplicative LoRA weights are aligned with base networks, once permutation symmetry is eliminated, as we prove in Corollary 2.2 in the Appendix.
Despite the application of the asymmetric mask, additive LoRA weights does not exhibit good linear mode connectivity as mLoRA does. We hypothesize this is related to the internal mechanisms of neural fields. Neural fields synthesize signals through iterative composition: linear layers combine basic signal components, while activation functions generate harmonics yuce2022structured. When neural fields are optimized on individual instances, their channels become highly entangled. However, when trained on multiple instances in a generalizable setting, the network learns transferable features and exhibits greater disentanglement radford2021learning. This observation provides additional motivation for fine-tuning a base network trained in the generative regime. Additive LoRA reintroduces entanglement by mixing features across channels, whereas multiplicative LoRA preserves the channel structure through aligned feature scaling, avoiding additional entanglement.
@@ 4.3 Generation via Diffusion Models
ShapeNet - Airplane
HyperDiffusion 2.39 43.6% 78.2% 0.027 0.009 0.122
MLP-Asym 2.80 44.8% 80.9% 0.041 0.018 0.254
LoRA 116.4 3.1% 99.9% 1.553 0.669 7.163
LoRA-Asym 270.6 2.0% 100% 1.532 0.823 7.931
mLoRA 1.96 46.2% 70.5% 0.049 0.025 0.359
mLoRA-Asym 1.89 43.4% 71.9% 0.011 0.003 0.041
ShapeNet - Multi
HyperDiffusion 8.64 41.6% 78.3% 0.117 0.023 0.219
MLP-Asym 7.77 46.5% 74.0% 0.085 0.016 0.157
LoRA 152.9 10.9% 99.1% 1.014 0.319 2.501
LoRA-Asym 210.0 1.2% 100% 1.241 0.437 2.987
mLoRA 5.75 46.4% 61.9% 0.071 0.011 0.123
mLoRA-Asym 5.52 49.6% 58.6% 0.026 0.004 0.040
@@
Implementation. For standalone MLP weights, we use the same diffusion model architecture as HyperDiffusion erkocc2023hyperdiffusion. For LoRA weights, we employ our hierarchical LoRA layer encoder described in Section 3.5. For standalone MLP weights and standalone MLP weights with asymmetric masks, we adopt the initialization technique from erkocc2023hyperdiffusion, which first fits one instance and then uses its weights to initialize all other fittings. This can be viewed as fine tuning a small MLP fitted to one instance to other instances, enabling direct comparison with the state-of-the-art weight space diffusion method.
Evaluation Metrics. For both 2D and 3D data, we measure the difference between the distribution of the generated and reference samples. We compute the Fréchet distance (FD) frechet1957distance, as well as the Maximum Mean Discrepancy (MMD) calculated using a Gaussian RBF kernel (MMD-G) and a polynomial kernel (MMD-P), which has been shown to be more reliable and sample efficient than the Fréchet distance heusel2017gans; binkowski2018demystifying. These metrics operate on features extracted by deep learned models, as deep learned feature extractors project the data into semantically rich embedding spaces where distances better correlate with human perception of similarity. We employ these metrics in their mathematical form rather than using modality-specific implementations like FID heusel2017gans or KID binkowski2018demystifying, as this allows for consistent evaluation across different data modalities. For 2D images, we use CLIP radford2021learning as the feature extractor; for 3D shapes, we use a PointNet++ qi2017pointnet++. For 3D shapes, following erkocc2023hyperdiffusion; vahdat2022lion; luo2021diffusion; zhou20213d, we also report distance-based metrics: Minimum Matching Distance (mMD), Coverage (COV), and 1 Nearest Neighbor Accuracy (1-NNA). These metrics use the Chamfer Distance to measure shape similarity. Formal definitions are presented in the supplementary material.
@@ 4.4 Discriminative Tasks
Results. Table 4 reports the accuracy for the two classification methods and the Adjusted Rand Score for clustering across the six candidate representations. We also present t-SNE visualizations of the weight representations for the ten object categories in Figure 5. Each point encodes one instance, colored by its category. The classification and clustering results reveal a clear progression in semantic structure across different parameterizations. LoRA outperforms standalone MLPs, mLoRA outperforms LoRA, and mLoRA delivers the best results overall, achieving $90\%$ accuracy with a linear classifier. Given that mLoRA-Asym exhibits better weight space structure in terms of linear mode connectivity, thsi result indicate that the discriminative power of the weight representation is not directly linked to linear mode connectivity or permutation symmetry. The t-SNE visualizaitons show that all weight representations are able to capture some level of semantic structure, as same-category instances are positioned closely. However, only mulitplicative LoRA weights demonstrate clear class seperation. These results substantiate that multiplicative LoRA weight representations are capable of capturing semantic structure. Notably, recent work weights2weights dravid2024interpreting similarly observes that LoRA weights that fine-tune image diffusion model exhibit interpretable structure with semantic linear directions, echoing our observations in the context of neural fields. Our finding further challenges the popular view that neural network weights are uninterpretable, demonstrating that with appropriate parameterization, weight spaces can encode semantic information.
5 Limitation
First, our approach requires finetuning a base-model which is computationally more expensive than fitting small MLPs. This requirement prevents evaluation on datasets with hundreds of thousands of samples, constraining our experiments to datasets with thousands of instances. Future work could explore methods to eliminate this requirement, or develop more computationally efficient multiplicative LoRA adaptation procedures.
Second, we apply uniform parameter budgets across all layers in the neural field, treating each layer identically in terms of LoRA rank and capacity. However, different layers likely encode different levels of semantic abstraction, suggesting that adaptive parameter allocation could improve efficiency and representation quality.
@@ 6 Conclusion
We have demonstrated that neural network weights can serve as effective data representations when constrained through appropriate inductive biases. By adapting a pre-trained base model via multiplicative LoRA, we transform the chaotic parameter space into structured weight representations that exhibit semantic organization, enabling superior reconstruction quality, generation performance, and semantic structure. Remarkably, mLoRA-Asym weights converge to a linear mode during optimization, and this structured weight space geometry correlates strongly with generative performance in diffusion models. These findings challenge the view of weights as opaque byproducts of optimization and establish their viability as semantic representations for reconstruction, generation, and discriminative tasks.
Appendix
@@ 7.1 Permutation Symmetry in Additive LoRA
This permutation symmetry is analogous to the well known permutation symmetry in standard MLPs zhou2023permutation: reordering hidden neurons (along with their corresponding incoming and outgoing weights) does not change the function computed by the network. In LoRA, the hidden dimension is the rank $r$, and permuting this dimension induces the symmetry. Figure 6(a) illustrates this symmetry, showing how the low rank matrices can be viewed as encoder and decoder layers where intermediate neurons can be reordered.
@@ 7.2 Permutation Symmetry in Multiplicative LoRA
$\mathbf{B}=\begin{bmatrix}\mathbf{b}_{1}&\mathbf{b}_{2}&\cdots&\mathbf{b}_{r}\end{bmatrix},\quad\mathbf{A}=\begin{bmatrix}\mathbf{a}_{1}^{T}\\ \mathbf{a}_{2}^{T}\\ \vdots\\ \mathbf{a}_{r}^{T}\end{bmatrix}$ (12)
@@
Once permutation symmetry is eliminated, multiplicative LoRA weights are aligned with the channels in the base network, i.e. each element corresponds to a single feature channel.
@@ 8.1 Standalone MLP
As shown in Figure 7(a), The standalone MLP is a Fourier Feature tancik2020fourier layer $\boldsymbol{\alpha}_{1}=\sin(\omega_{0}\cdot(\mathbf{W}_{1}\mathbf{p}+\mathbf{b}_{1}))$ followed by $2$ linear layers.
@@
As noted by Erkoç et al. erkocc2023hyperdiffusion, an initialization trick is employed to ensure diffusion model generalization. Specifically, an MLP with weights $\boldsymbol{\phi}_{0}$ is fitted to one instance from the dataset, and used as a shared initialization for all the rest of the fittings $\boldsymbol{\iota}_{i}=\boldsymbol{\phi}_{0}$. In other words, other instances are fitted by fine-tuning the weights from the first instance. This trick is crucial to the performance of HyperDiffusion, and therefore we employ it in the FFHQ and ShapeNet airplane experiments. However, in the multi-category experiment, it does not make sense to initialize a chair fitting with weights from an airplane, therefore we do not use this trick and instead use a shared random initilization for all the fittings, like is done for the LoRA-based weight representations.
@@ 8.2 Base Model Architecture
The base model follows the modulated neural field architecture widely adopted in generative neural field works anokhin2021cips; karras2021alias; chan2020pigan, as illustrated in Figure 7(b). The architecture consists of a mapping network and a synthesis network. The mapping network is a multilayer perceptron that transforms a latent code $\mathbf{z}\in\mathbb{R}^{d_{z}}$ into intermediate style vectors $\{\mathbf{s}_{l}\}_{l=1}^{L}$, where $L$ is the number of synthesis layers. For 2D FFHQ, we use $d_{z}=256$ and an 8 layer mapping network with hidden dimension 256. For 3D ShapeNet, we use $d_{z}=128$ and a 4 layer mapping network with hidden dimension 256.

### v2

@@ 3.5 Diffusion Model on Weight Representations
We parameterize the reverse process using a diffusion transformer (DiT) [31] that predicts the noise added to the weights. For standalone MLP weights, we adopt the architecture from [9, 30]. For LoRA weights, we design a hierarchical LoRA layer encoder module that respects the structural properties of low-rank weight matrices, shown in Figure 2. Each layer $l$ is processed as follows. First, we treat each vector pair $(\mathbf{a}_{l}^{(i)},\mathbf{b}_{l}^{(i)})$ as a token, where $\mathbf{a}_{l}^{(i)}$ and $\mathbf{b}_{l}^{(i)}$ are the $i$-th columns and rows of matrices $\mathbf{A}^{(l)}$ and $\mathbf{B}^{(l)}$, respectively. Vector-level positional encodings are then applied to capture the rank dimension index. A multi-head attention module with $r$ attention heads encodes the interactions among the $r$ vector pairs within the layer, allowing the model to learn dependencies between different rank components. Finally, layer-level positional encodings are applied to the aggregated layer representation before feeding into the main transformer.
@@ 4 Experiments
Datasets. We evaluate on two datasets. For 2D data, we use FFHQ [21], which contains high quality face images. We evaluate at the resolution of $128\times 128$, although modest compared to current state-of-the-art in image generation, this resolution is significantly higher than what previous weight-space methods [46, 45] have evaluated on and therefore more challenging. For 3D data, we use ShapeNet [4], focusing on two settings: A single-category model trained on airplanes, and a multi-category model trained on ten object categories including airplanes, chairs, tables, and other common objects.
@@ 4.1 Reconstruction
Results. As shown in Table 1, LoRA, mLoRA and mLoRA-Asym achieve better reconstruction quality while maintaining a compact parameter count, compared to the MLP-based weight representations. We attribute this to the inductive bias provided by the base network, which captures transferable features across instances, and fine tuning effectively leverages these shared representations with minimal adaptation parameters. Multiplicative LoRA perform better than its additive counterpart, echoing the wide application of multiplicative modulation in generative neural fields [1, 3, 20]. Surprisingly, mLoRA-Asym outperforms mLoRA despite having certain parameters frozen by the asymmetric mask. Our hypothesis is that the masking further reduces parameter entanglement and thereby improves the reconstruction accuracy. Also notably, LoRA-Asym performs poorly on FFHQ, likely due to increased parameter entanglement caused by large variance $\kappa=6$ (empirically determined) of the frozen weights, as discussed in Section  3.4.
@@ 4.2 Weight Space Structure Analysis
Second, to examine linear mode connectivity [11], we measure the barrier height by evaluating reconstruction quality at the midpoint of the linear interpolation path. Specifically, we compute the Chamfer Distance between the ground truth mesh vertices $\mathbf{v}_{\text{gt}}$ and the vertices $\mathbf{v}_{\text{avg}}$ extracted via marching cubes from the averaged weights, i.e.,
@@
Although not always improving reconstruction quality, the asymmetric mask improves both weight similarity and linear mode connectivity across all parameterizations. Notably, mLoRA-Asym exhibits exceptional behavior: weight similarity remains very high and the barrier remains very low even with very different initializations. This suggests that mLoRA-Asym weights converge to a linear mode. The reason is likely that multiplicative LoRA weights are aligned with base networks, once permutation symmetry is eliminated, as we prove in Corollary 2.2 in the Supplementary Materials.
Despite the application of the asymmetric mask, the use of additive LoRA weights does not exhibit as good a linear mode connectivity as mLoRA does. We hypothesize that this is related to the internal mechanisms of neural fields. Neural fields synthesize signals through iterative composition: linear layers combine basic signal components, while activation functions generate harmonics [43]. When neural fields are optimized on individual instances, their channels become highly entangled. However, when trained on multiple instances in a generalizable setting, the network learns transferable features and exhibits greater disentanglement [33]. This observation provides additional motivation for fine-tuning a base network trained in the generative regime. Additive LoRA reintroduces entanglement by mixing features across channels, whereas multiplicative LoRA preserves the channel structure through aligned feature scaling, avoiding additional entanglement.
@@ 4.3 Generation via Diffusion Models
ShapeNet - Airplane ShapeNet - Multi
mMD $\downarrow$ COV $\uparrow$ 1-NNA $\downarrow$ FD $\downarrow$ MMD-G $\downarrow$ MMD-P $\downarrow$ mMD $\downarrow$ COV $\uparrow$ 1-NNA $\downarrow$ FD $\downarrow$ MMD-G $\downarrow$ MMD-P $\downarrow$
HyperDiffusion [9] 2.39 43.6% 78.2% 0.027 0.009 0.122 8.64 41.6% 78.3% 0.117 0.023 0.219
MLP-Asym 2.80 44.8% 80.9% 0.041 0.018 0.254 7.77 46.5% 74.0% 0.085 0.016 0.157
LoRA 116.4 3.1% 99.9% 1.553 0.669 7.163 152.9 10.9% 99.1% 1.014 0.319 2.501
LoRA-Asym 270.6 2.0% 100% 1.532 0.823 7.931 210.0 1.2% 100% 1.241 0.437 2.987
mLoRA 1.96 46.2% 70.5% 0.049 0.025 0.359 5.75 46.4% 61.9% 0.071 0.011 0.123
mLoRA-Asym 1.89 43.4% 71.9% 0.011 0.003 0.041 5.52 49.6% 58.6% 0.026 0.004 0.040
@@
Implementation. For standalone MLP weights, we use the same diffusion model architecture as HyperDiffusion [9]. For LoRA weights, we employ our hierarchical LoRA layer encoder described in Section 3.5. For standalone MLP weights and standalone MLP weights with asymmetric masks, we adopt the initialization technique from [9], which first fits one instance and then uses its weights to initialize all other fittings. This can be viewed as fine tuning a small MLP fitted to one instance to other instances, enabling direct comparison with the state-of-the-art weight space diffusion method.
Evaluation Metrics. For both 2D and 3D data, we measure the difference between the distribution of the generated and reference samples. We compute the Fréchet distance (FD) [12], as well as the Maximum Mean Discrepancy (MMD) calculated using a Gaussian RBF kernel (MMD-G) and a polynomial kernel (MMD-P), which has been shown to be more reliable and sample efficient than the Fréchet distance [16, 2]. These metrics operate on features extracted by deep learned models, as deep learned feature extractors project the data into semantically rich embedding spaces where distances better correlate with human perception of similarity. We employ these metrics in their mathematical form rather than using modality-specific implementations like FID [16] or KID [2], as this allows for consistent evaluation across different data modalities. For 2D images, we use CLIP [33] as the feature extractor; for 3D shapes, we use a PointNet++ [32]. For 3D shapes, following [9, 37, 26, 47], we also report distance-based metrics: Minimum Matching Distance (mMD), Coverage (COV), and 1 Nearest Neighbor Accuracy (1-NNA). These metrics use the Chamfer Distance to measure shape similarity. Formal definitions are presented in the supplementary material.
@@ 4.4 Discriminative Tasks
Results. Table 4 reports the accuracy for the two classification methods and the Adjusted Rand Score for clustering across the six candidate representations. We also present t-SNE visualizations of the weight representations for the ten object categories in Figure 5. Each point encodes one instance, colored by its category. The classification and clustering results reveal a clear progression in semantic structure across different parameterizations. LoRA outperforms standalone MLPs, mLoRA outperforms LoRA, and mLoRA delivers the best results overall, achieving $90\%$ accuracy with a linear classifier. Given that mLoRA-Asym exhibits better weight space structure in terms of linear mode connectivity, this result indicates that the discriminative power of the weight representation is not directly linked to linear mode connectivity or permutation symmetry. The t-SNE visualizations show that all weight representations are able to capture some level of semantic structure, as same-category instances are positioned closely. However, only multiplicative LoRA weights demonstrate clear class separation. These results substantiate that multiplicative LoRA weight representations are capable of capturing semantic structure.
@@ 5 Conclusion
We have demonstrated that independently optimized neural network weights can serve as effective data representations when constrained through appropriate inductive biases. By adapting a pre-trained base model via multiplicative LoRA, we transform the chaotic parameter space into structured weight representations that exhibit semantic organization, enabling superior reconstruction quality, generation performance, and semantic structure. Remarkably, mLoRA-Asym weights converge to a linear mode during optimization, and this structured weight space geometry correlates strongly with generative performance in diffusion models. These findings challenge the view of weights as opaque byproducts of optimization and establish their viability as semantic representations for reconstruction, generation, and discriminative tasks.
Supplementary Material
This supplementary material provides additional theoretical analysis, implementation details, and experimental results to support the main paper. Section 6 presents formal proofs demonstrating permutation symmetry in both additive and multiplicative LoRA parameterizations. Section 7 discusses related works in implicit neural representations. Section 8 provides comprehensive implementation details for the standalone MLP architecture, base model architecture and training, dataset preparation, and the complete training pipeline. Section 9 formally defines the evaluation metrics used throughout our experiments. Section 10 presents an ablation study examining the effectiveness of the hierarchical LoRA layer encoder. Section 11 presents weight space interpolation experiments. Section 12 provides additional qualitative generation results on FFHQ and ShapeNet datasets. Finally, Section 13 discusses limitations and future research directions.
@@ 6.1 Permutation Symmetry in Additive LoRA
This permutation symmetry is analogous to the well known permutation symmetry in standard MLPs [45]: reordering hidden neurons (along with their corresponding incoming and outgoing weights) does not change the function computed by the network. In LoRA, the hidden dimension is the rank $r$, and permuting this dimension induces the symmetry. Figure 6(a) illustrates this symmetry, showing how the low rank matrices can be viewed as encoder and decoder layers where intermediate neurons can be reordered.
@@ 6.2 Permutation Symmetry in Multiplicative LoRA
$\mathbf{B}=\begin{bmatrix}\mathbf{b}_{1}&\mathbf{b}_{2}&\cdots&\mathbf{b}_{r}\end{bmatrix},\qquad\mathbf{A}=\begin{bmatrix}\mathbf{a}_{1}^{T}\\ \mathbf{a}_{2}^{T}\\ \vdots\\ \mathbf{a}_{r}^{T}\end{bmatrix}$ (12)
@@
Once permutation symmetry is eliminated, multiplicative LoRA weights are completely aligned with the channels in the base network.
@@
7 Related Works - Implicit Neural Representations
Implicit Neural Representations (INRs), also known as neural fields, are continuous functions parameterized by neural networks that map coordinates to signal values. INRs represent signals as continuous functions $\Phi:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$, where a neural network maps $n$-dimensional coordinates to $m$-dimensional quantities. This paradigm enables resolution-independent and modality agnostic representations of complex signals [41, 10], employing identical network architectures across diverse modalities including 1D audio, 2D images, 3D shapes, and even 4D spatiotemporal data.
Beyond single-instance fitting, generalizable INRs have been proposed to learn priors across datasets through approaches based on autoencoders [29], generative adversarial networks (GANs) [1, 20, 3] and shared layers [38]. The GAN-based works could be seen as extensions of the StyleGAN [21] paradigm into the realm of neural fields. They generate different instances by modulating an MLP trunk, which we use as the base network for fine-tuning.
Because INRs parameterize data as neural network functions, the weights offer a direct pathway to data representation. This perspective has practical applications in compression, with methods [7, 14] demonstrating competitive compression ratios by storing quantized network parameters instead of raw data. However, whether the collection of weights could encode semantic structure remains an open question. Another line of work employs hypernetworks [22] and transformers [5] to predict INR weights from input data via learned mappings as a way of data generation. Dupont et al. [8] propose functa, which meta-learns a shared SIREN base network and represents each data point as a low-dimensional shift modulation vector for downstream tasks including generation and classification.
In contrast to all the above approaches, we investigate whether independently optimized weights can directly serve as meaningful representations. HyperDiffusion [9] trains a diffusion transformer to generate neural field weights as a means of synthesizing 3D shapes and 4D animated shapes. Our work builds on this to inspect factors that affect weight space generation performance and to explore semantic structures in neural field weights.
@@ 8.1 Standalone MLP
As shown in Figure 7(a), The standalone MLP is a Fourier Feature [36] layer $\boldsymbol{\alpha}_{1}=\sin(\omega_{0}\cdot(\mathbf{W}_{1}\mathbf{p}+\mathbf{b}_{1}))$ followed by $2$ linear layers.
@@
As noted by Erkoç et al. [9], an initialization trick is employed to ensure diffusion model generalization. Specifically, an MLP with weights $\boldsymbol{\phi}_{0}$ is fitted to one instance from the dataset, and used as a shared initialization for all the rest of the fittings $\boldsymbol{\iota}_{i}=\boldsymbol{\phi}_{0}$. In other words, other instances are fitted by fine-tuning the weights from the first instance. This trick is crucial to the performance of HyperDiffusion, and therefore we employ it in the FFHQ and ShapeNet airplane experiments. However, in the multi-category experiment, it does not make sense to initialize a chair fitting with weights from an airplane, therefore we do not use this trick and instead use a shared random initilization for all the fittings, like is done for the LoRA-based weight representations.
@@ 8.2 Base Model Architecture
The base model follows the modulated neural field architecture widely adopted in generative neural field works [1, 20, 3], as illustrated in Figure 7(b). The architecture consists of a mapping network and a synthesis network. The mapping network is a multilayer perceptron that transforms a latent code $\mathbf{z}\in\mathbb{R}^{d_{z}}$ into intermediate style vectors $\{\mathbf{s}_{l}\}_{l=1}^{L}$, where $L$ is the number of synthesis layers. For 2D FFHQ, we use $d_{z}=256$ and an 8 layer mapping network with hidden dimension 256. For 3D ShapeNet, we use $d_{z}=128$ and a 4 layer mapping network with hidden dimension 256.